OpenStudy (anonymous):

my problem is 9(3x+9)=14-(x+3) so far i have worked it out but i am now stuck here is what i have done 9(3x+9)=14-(x+3) 27x+9=14-x+3 27x+9=17-x 28x+9=17-x+x 28x+9=17 28x+9-9=18-9 28x/28=9/28 i know that 28x over 28 the leves just x=9/28 but i need to simplify it and i dont know how

OpenStudy (anonymous):

made a mistake in line one

OpenStudy (anonymous):

or rather line two

OpenStudy (anonymous):

rhs is \[14-(x+3)=14-x-3\]

OpenStudy (anonymous):

you did multiply 9 to the other 9

OpenStudy (anonymous):

also need distributive law on lhs as well \[9(3x+9)=27x+81\]

OpenStudy (anonymous):

ya lol we rainbow multiplication :P or well thats what i call it...

OpenStudy (anonymous):

step one should be to write \[27x+81=14-x-3\]

OpenStudy (anonymous):

rainbow? hurray for new math!

OpenStudy (anonymous):

it is called the "distributive law" or maybe the "distributive property"

OpenStudy (anonymous):

The answer is ( x = -3)

OpenStudy (anonymous):

we can finish easy \[27x+81=11-x\] \[27x+x=11-81\] \[28x=-70\] \[x=\frac{-70}{28}=-\frac{5}{2}\]

OpenStudy (anonymous):

oops did i mess up?

OpenStudy (anonymous):

thank you satellite you got it right thanks

OpenStudy (anonymous):

no i am sticking to my answer of \[-\frac{5}{2}\]

OpenStudy (anonymous):

you need to factor the 9(3x +9) = 14-(x + 3) 27x +81 = 14(-x - 3) +14x -81 +14x -81 41x = -123 x = -3

OpenStudy (anonymous):

put the answer i just gave you with one equals sign per line and you will understand

OpenStudy (anonymous):

would u like me to send you this problem worded out in an email?

OpenStudy (anonymous):

how do i post a new question i cant delete what in the ask box

OpenStudy (anonymous):

You don't have to delete. just go to left and post where says 'ask a question.'

OpenStudy (anonymous):

Shandi.bilbrey it's 14 minus (x+3) not 14 times minus (x+3)

OpenStudy (anonymous):

hat i was saying is you needed to distribute the negitive

OpenStudy (anonymous):

yes but once you've distributed the negative sign you don't need the brackets What satelitte73 did was correct